OpenStudy (anonymous):

Help here please: Do some research to find an investment, like a GIC or a Canada Savings Bond

OpenStudy (experimentx):

seems the same problem. do you know the equation for regular deposit compound interest problem,

OpenStudy (experimentx):

let's google ...

OpenStudy (phi):

Here are the rates for Canada Savings Bond, http://www.csb.gc.ca/about/rates/

OpenStudy (experimentx):

you could use the calc given by directrix on earlier question to solve it.

OpenStudy (experimentx):

i got one http://www.ajdesigner.com/phpinterest/interest_regular_deposits_p.php

OpenStudy (phi):

Say you picked Canada Premium Bonds - Outstanding Series The outstanding CPB series listed below will carry the following rates: Year 1: 1.00% You can use this formula to find the Future Value of $1 200 each year ($100 per month) for five years: FV= Monthly_Payment \( \cdot \frac{(1+i)^n - 1}{i} \) In the formula, i is the \(monthly\) interest rate = 1%/12 = 0.01/12= \(\frac{1}{1200}\) n is the time in months. So 5 years is 60 months The monthly payment is $100.

OpenStudy (phi):

Now we plug in the numbers into the formula FV= 100 \( \frac{(1 + \frac{1}{1200})^{60} -1)}{\frac{1}{1200}} \) You need a calculator to figure this out.

OpenStudy (phi):

First, what do you get evaluating the FV formula up above?

OpenStudy (phi):

6149.90 you put in 100 per month for 60 months (5 years)= 6000.00 so the interest would be 149.90

OpenStudy (phi):

Now you can do What if you contributed $2 400 each year? which is just 200 per month, using the same FV formula

OpenStudy (anonymous):

12299.80. So it would be worth investing my money right?

OpenStudy (phi):

Is it worth it? That is up to you.